# a-DCF: an architecture agnostic metric with application to spoofing-robust speaker verification

### v1

@@ 1 Introduction
The t-DCF, however, is somewhat restrictive in terms of supported architectures. The speaker detector is used as a gate to the spoof detector,222There are plenty of other possible combination architectures. while an AND decision logic is used to combine their respective classification decisions; given a target trial, the spoof detector should indicate that an input utterance is bonafide AND the speaker detector should indicate that the input utterance corresponds to the claimed identity. The t-DCF requires the computation of separate speaker and spoof detection scores and cannot be applied to the evaluation of alternative approaches which, for example, might produce only a single score. Such approaches have been reported in the past and continue to emerge [14, 15, 6, 16, 17, 18, 19]. Their incompatibility with the t-DCF metric, has hence stimulated the development of alternative metrics, all forms of equal error rate (EER) estimates [14]. Different metrics are hence employed for the evaluation of different architectural approaches to the evaluation of spoofing-robust ASV solutions, despite them all sharing exactly the same goal. Use of different metrics then complicates the benchmarking of competing solutions.
@@ 2 Related work
The tandem-detection cost function (t-DCF) [12] was developed for this purpose. Again, there were several advantages, not least because it provided an alternative to the EER metric. As a result of the assumed conditional independence of speaker and spoof detection decisions, the t-DCF also promotes the independent development of spoof detectors.333The conditional independence assumption of the two sub-systems even enables them to be developed using different datasets, yet they can still be evaluated in combination. Decisions produced by a spoof detector can be combined with those made by a speaker detector to provide a measure of spoofing-robust ASV performance. At the same time, use of a DCF-derived metric also helped bridge the gap between the relatively young spoof detection community and the mature speaker detection community which had embraced the DCF metric [20] decades earlier.
@@
The ensemble of metrics provides insights into classifier performance under different trial combinations, as well as the impacts of spoofing and spoof detection upon ASV (just like the t-DCF). Even so, there are three metrics (instead of one), and they are all EER estimates (instead of DCF estimates). The shortcomings of any EER-based metric are well-known. For ASV applications, one generally prioritises either a low miss rate or a low false alarm rate, which conflicts with the notion of equal with the EER. Unlike DCF-based metrics, the EER is not customizable or an optimization target for different applications. The SASV-EER metric which pools non-targets and spoofing attacks among the negative class has even further issues. The classic EER, when used to measure discrimination between any pair of two classes is not dependent on the empirical class priors. As soon as the third class (spoofing attacks) is pooled with non-targets, however, this property no longer holds [13, Section 4.4]; the SASV-EER metric becomes a function of the empirical class priors. Hence, the metric itself depends on the evaluation data.
@@ 3.1 General form
where is a vector of ones which acts to sum the columns of the matrix resulting from the terms to the right, and where is the Hadamard or entry-wise product.
@@ 3.2 Spoofing-robust speaker verification
When subjected to spoofing attacks, the non-target input class becomes a union of bonafide (zero-effort impostor) non-target trials and spoofed (target) trials. For brevity, we refer to them simply as non-target and spoofed trials respectively. Different priors and costs can be assigned to each. Even so, the decision is still binary, with inputs still being labeled as either targets (positive class) or non-targets (negative class). Identical to the traditional ASV scenario (without spoofing attacks), there hence remains a single cost for missed target detections. However, since there are two different negative class priors, one for non-targets and one for spoofs,444Note that positive and negative class assignments are arbitrary [12]. In keeping with the literature, here we assume the assingment of bonafide target trials to the positive class and everything else to the negative class. there are likewise two (possibly different)555From an applications perspective, the mistaking of non-targets for targets when caused by an unintentional impostor may not have the same ramifications (cost) of when a spoof mistaken for a target is caused by a fraudster. In the spirit of ‘an error is an error, no matter the cause’, for all experiments reported later in the paper, we use the same costs for each type of false alarm. false alarm costs. Classifier conditional probabilities can also be estimated in the usual way for a specified operating point or threshold .
@@ 3.3 Normalisation
While the normalised a-DCF function in (8) can exceed 1, which implies that it performs worse than a random system which either accepts or rejects every trial, the minimum is always bounded between 0 and 1. All detection costs reported later in the paper are normalised minimums computed according to (9). For simplicity, they are denoted simply as a-DCF.
3.4 a-DCF as generalisation of NIST DCF and a special case of t-DCF
It would be helpful to discuss the relation between a-DCF and other DCFs reported in [20] and [12]. The comparison is shown in Table 2. The original DCF [20] in (1), as endorsed by NIST through the speaker recognition evaluation (SRE) series, is designed for the benchmarking of binary classifiers which produce a single detection score — such as standalone speaker or spoof detection systems. The NIST DCF can be seen as a degenerate case of the a-DCF with (no spoofing attacks).
@@ 4.2 Systems
We consider three types of systems, cascade, jointly optimised, and single-model, to demonstrate the architecture-agnostic of the a-DCF metric.101010Note that although we categorise systems in Table 5 into three categories, it does not mean that all systems should fall into one of the three categories. ‘Jointly optimised’ can be defined differently under diverse circumstances, where in our case, we use the term to call systems that jointly optimise pre-trained ASV and CM systems to derive the final score for each input. For the cascade (decision-level tandem) systems, RawNet3 [26] is used as an ASV model. As a CM system, AASIST [27], AASIST-L [27] are used and several cascade systems are composed. In addition, we consider systems submitted to the SASV 2022 challenge [14, 29, 30, 31] which explored diverse methodologies including score-level and embedding-level combinations. For the single-model systems, we employ various versions from [32], comprising MFA-Conformer [33] and SKA-TDNN [34] models, where the training data varies throughout the devised multi-stage training.111111S1, S2, S3, and S4 represent MFA-Conformer index 7 and SKA-TDNN indices 3, 7, and 11, respectively, from the model indices shown in https://github.com/sasv-challenge/SASV2_Baseline. To avoid confusion regarding system details and to focus on comparing metric measurements, all systems are denoted simply with the alphanumeric identifiers in Table 5.
@@ 7 Acknowledgement
We appreciate Dr. Shinji Watanabe for the valuable discussions.
The work has been partially supported by the Academy of Finland (Decision No. 349605, project “SPEECHFAKES”) and by the Agence Nationale de la Recherche (ANR) in France (BRUEL project, No. ANR-22-CE39-0009).
Experiments of this work used the Bridges2 system at PSC and Delta system at NCSA through allocations CIS210014 and IRI120008P from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.

### v2

@@ 1 Introduction
The t-DCF, however, is somewhat restrictive in terms of supported architectures. The speaker detector is used as a gate to the spoof detector,222There are plenty of other possible combination architectures. while an AND decision logic is used to combine their respective classification decisions; given a target trial, the spoof detector should indicate that an input utterance is bonafide AND the speaker detector should indicate that the input utterance corresponds to the claimed identity. The t-DCF requires the computation of separate speaker and spoof detection scores and cannot be applied to the evaluation of alternative approaches which, for example, might produce only a single score. Such approaches have been reported in the past and continue to emerge [14, 15, 6, 16, 17, 18, 19]. Their incompatibility with the t-DCF metric, has hence stimulated the development of alternative metrics, all forms of equal error rate (EER) estimates [14]. Different metrics are hence employed for the evaluation of different architectural approaches to the evaluation of spoofing-robust ASV solutions, despite them all sharing exactly the same goal. Use of different metrics then complicates the benchmarking of competing solutions. Furthermore, while an evaluation metric introduced in [20] can be adapted to incorporate the aforementioned condition, no existing research explicitly defines this problem to date.
@@ 2 Related work
The tandem-detection cost function (t-DCF) [12] was developed for this purpose. Again, there were several advantages, not least because it provided an alternative to the EER metric. As a result of the assumed conditional independence of speaker and spoof detection decisions, the t-DCF also promotes the independent development of spoof detectors.333In principle, the conditional independence assumption of the two sub-systems even enables them to be developed using different datasets, yet they can still be evaluated in combination; in practice, this is not recommended, as it introduces an unnecessary data variation factor. Decisions produced by a spoof detector can be combined with those made by a speaker detector to provide a measure of spoofing-robust ASV performance. At the same time, use of a DCF-derived metric also helped bridge the gap between the relatively young spoof detection community and the mature speaker detection community which had embraced the DCF metric [21] decades earlier.
@@
The ensemble of metrics provides insights into classifier performance under different trial combinations, as well as the impacts of spoofing and spoof detection upon ASV (just like the t-DCF). Even so, there are three metrics (instead of one), and they are all EER estimates (instead of DCF estimates). The shortcomings of any EER-based metric are well-known [3, 25, 13]. For ASV applications, one generally prioritises either a low miss rate or a low false alarm rate, which conflicts with the notion of equal with the EER. Unlike DCF-based metrics, the EER is not customizable or an optimization target for different applications. The SASV-EER metric which pools non-targets and spoofing attacks among the negative class has even further issues. The classic EER, when used to measure discrimination between any pair of two classes is not dependent on the empirical class priors. As soon as the third class (spoofing attacks) is pooled with non-targets, however, this property no longer holds [13, Section 4.4]; the SASV-EER metric becomes a function of the empirical class priors. Hence, the metric itself depends on the evaluation data.
@@ 3.1 General form
where is a vector of ones which acts to sum the columns of the matrix resulting from the terms to the right, and where is the Hadamard or entry-wise product. Though expressed using different formalism, (4) coincidences with [26, Eq. (4)].
@@ 3.2 Spoofing-robust speaker verification
When subjected to spoofing attacks, the negative input class becomes a union of bonafide non-target (zero-effort impostor) trials and spoofed target trials. For brevity, we refer to them simply as non-target and spoofed trials respectively. Different priors and costs can be assigned to each. Even so, the decision is still binary, with inputs still being labeled as either targets (positive class) or non-targets (negative class). Identical to the traditional ASV scenario (without spoofing attacks), there hence remains a single cost for missed target detections. However, since there are two different negative class priors, one for non-targets and one for spoofs,444Note that positive and negative class assignments are arbitrary [12]. In keeping with the literature, here we assume the assignment of bonafide target trials to the positive class and everything else to the negative class. there are likewise two (possibly different)555From an applications perspective, the mistaking of non-targets for targets when caused by an unintentional impostor may not have the same ramifications (cost) of when a spoof mistaken for a target is caused by a fraudster. In the spirit of ‘an error is an error, no matter the cause’, for all experiments reported later in the paper, we use the same costs for each type of false alarm. false alarm costs. Classifier conditional probabilities can also be estimated in the usual way for a specified operating point or threshold .
@@ 3.3 Normalisation
While the normalised a-DCF function in (8) can exceed 1, which implies that it performs worse than a non-informative system which either accepts or rejects every trial, the minimum is always bounded between 0 and 1. All detection costs reported later in the paper are normalised minimums computed according to (9). For simplicity, they are denoted simply as a-DCF.
3.4 Relation to NIST DCF and t-DCF
It would be helpful to discuss the relation between a-DCF and other DCFs reported in [21, 20, 12]. The comparison is shown in Table 2. The original DCF [21] in (1), as endorsed by NIST through the speaker recognition evaluation (SRE) series, is designed for the benchmarking of binary classifiers which produce a single detection score — such as standalone speaker or spoof detection systems. The NIST DCF can be seen as a degenerate case of the a-DCF with (no spoofing attacks). Interestingly, the NIST 2012 SRE campaign [20] used a 3-term DCF similar to (3.2), though the two FA rate terms are those of known/unknown nontargets (as opposed to nontargets and spoofs as in our work). Though the 2012 DCF and -DCF are in the same function class of evaluation metrics (i.e. differ only by the choice of costs and priors), they are motivated (and used) from entirely different perspectives.
@@ 4.2 Systems
We consider three types of systems: cascade, jointly optimised, and single-model, to demonstrate the architecture-agnostic of the a-DCF metric.101010Note that although we categorise systems in Table 5 into three categories, it does not mean that all systems should fall into one of the three categories. ‘Jointly optimised’ can be defined differently under diverse circumstances, where in our case, we use the term to call systems that jointly optimise pre-trained ASV and CM systems to derive the final score for each input. For the cascade (decision-level tandem) systems, RawNet3 [29] is used as an ASV model. As a CM system, AASIST [30], AASIST-L [30] are used and several cascade systems are composed. In addition, we consider systems submitted to the SASV 2022 challenge [14, 32, 33, 34] which explored diverse methodologies including score-level and embedding-level combinations. For the single-model systems, we employ various versions from [35], comprising MFA-Conformer [36] and SKA-TDNN [37] models, where the training data varies throughout the devised multi-stage training.111111S1, S2, S3, and S4 represent MFA-Conformer index 7 and SKA-TDNN indices 3, 7, and 11, respectively, from the model indices shown in https://github.com/sasv-challenge/SASV2_Baseline. To avoid confusion regarding system details and to focus on comparing metric measurements, all systems are denoted simply with the alphanumeric identifiers in Table 5.
@@ 7 Acknowledgement
We appreciate Dr. Shinji Watanabe for the valuable discussions. The work has been partially supported by the Academy of Finland (Decision No. 349605, project “SPEECHFAKES”) and by the Agence Nationale de la Recherche (ANR) in France (BRUEL project, No. ANR-22-CE39-0009). Experiments of this work used the Bridges2 system at PSC and Delta system at NCSA through allocations CIS210014 and IRI120008P from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.